OpenStudy (anonymous):

The regular price of an item is $84. The item is on sale with a discount of $12. What is the discount rate? A.0.143% B.1.43% C.14.3% D.143% -medal will be given

OpenStudy (anonymous):

-if u know or are fans of @Angelsoft can u plz add me . I was having trouble logging in with that account so I made this one can the following add me ?! @baconbaby1334 @kakes1967 @taylor115820 @dianolove Please and thanks for ur understanding!

OpenStudy (kropot72):

The discount rate is found as follows: \[\large \frac{12}{84}\times\frac{100}{1}=(you\ can\ calculate)\ percent\]

OpenStudy (anonymous):

ok thx very much @kropot72

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Tp now what am I going to call you